### v1

# REPS: Recycling Entropies for Packet Spraying to Adaptively Explore Paths and Mitigate Failures

Abstract.
Most existing datacenter transport protocols rely on in-order packet delivery, a design choice rooted in legacy systems and simplicity. However, advancements in technology, such as RDMA, have made it feasible to relax this requirement, allowing for more effective use of modern datacenter topologies like FatTree and Dragonfly. The rise of AI/ML workloads underscores the necessity for enhanced link utilization, a challenge for single-path load balancers due to issues like ECMP collisions.
In this paper, we introduce REPS, a novel per-packet traffic load-balancing algorithm that integrates seamlessly with existing congestion control mechanisms. REPS reroutes packets around congested hotspots and unreliable or failing links with remarkable simplicity and minimal state requirements.
Our evaluation demonstrates that REPS significantly outperforms traditional packet spraying and other state-of-the-art solutions in contemporary datacenter networks, offering substantial improvements in performance and link utilization.
1. Introduction
With the recent explosion in AI/ML networks, the battle for better next-generation transport protocols has heated up with the recent release of Poseidon (Wang et al., 2023) by Google and the incoming release of Ultra Ethernet (Consortium, 2024) by major tech players. Traditionally, AI/ML clusters have been using Infiniband and more recently Ethernet based solutions such as RoCEv2 (Association, 2024) to reduce cost and offer an easier to deploy alternative.
However, existing solutions struggle to keep up with the increasingly demanding AI/ML workload where the expected load is much higher than what encountered in cloud workloads. This increased load requirement showcases even more one of the main shortcomings of existing solutions: the routing algorithm. This is because, most of the times, the expectation is that each message will be routed using a single path to avoid out-of-order packets and reduce complexity of the load balancer. However, this results in ECMP collisions which reduce the overall performance of the whole transport layer due to queues building up and re-transmissions or PFC kicking in for lossless networks.
Moreover, a number of recent industry papers have highlighted the drastic effect of failures in both training times and economic costs (Qian et al., 2024). This highlights the importance of designing a transport layer with a load balancing scheme capable of adapting almost instantaneously to both network failures and asymmetries in the networks.
For this reasons, we propose REPS, a simple load balancing scheme based on adaptive packet spraying. We design REPS in order to work with any congestion control scheme and transport layer as long as it is able to handle out-of-order packets which is what next generation protocols are pushing for given the inherit performance advantage.
REPS does not require any specific hardware support from switches beside ECMP which is commonly present in modern switches. If available, REPS can take advantage of packet trimming to further improve its accuracy but it is also capable of running without that.
1.1. Background and Motivation
Modern datacenters tend to use Equal Cost Multi-Path (ECMP) (Hopps, 2009) as their default routing scheme for their switches. ECMP uses the 5-tuple header of each data packet: the protocol number, the IP addresses and the TCP or UDP source and destination port numbers. It uses such information to determine the next link to take by using a hash function. Packets that belong to the same connection will normally be routed using the same exact path from source to destination, which is the case for RoCEv2 for instance. This helps keeping the design trivial since, ignoring failures, it is extremely unlikely to receive out-of-order packets at the receiver. On the other hand, if two different connections get hashed to the same link, there will be a ECMP collision resulting in queues building up and in more extreme cases re-transmissions.
To mitigate such issues certain protocols have proposed splitting each flow into sub-flow and then routing each sub-flow individually, usually by artificially changing the source port field. This is the case of Flowlet Switching (Vanini et al., 2017), Flowcell (He et al., 2015) or PLB (Kabbani et al., 2014; Qureshi et al., 2022) that attempt to provide a trade-off between low amount of re-ordering and better performance.
However, if we can design a new transport layer that can relax such out-of-orderliness requirement, then we can theoretically do such operation at the individual packet by choosing a random source port field for each packet. This is called packet spraying (Dixit et al., 2013) and by its own, it is able to offer a significant performance improvement over single path algorithms or intermediate solutions such as Flowlet Switching.
Finally, we observe how we can take this one step further by adaptively spraying packets over paths. This, surprisingly, allows to achieve better results even with a perfectly symmetric network and an even bigger advantage when encountering failures.
1.2. Contribution
- REPS:
A powerful and lightweight adaptive load balancing scheme called Recycled Entropy Packet Spraying (REPS, Sec.3). REPS achieves near-optimal performance at high bandwidth utilization and efficiently routes around path failures, congestion hotspots, and network asymmetries. Unlike more complex in-network load balancing solutions(De Sensi et al., 2019; Rocher-González et al., 2022), REPS does not require any switch support beyond ECMP. REPS provides a significant improvement over traditional packet spraying, both in perfectly symmetric networks and in the presence of asymmetries, whether due to differing topologies or potential failures. We demonstrate that REPS performs well regardless of the congestion control scheme used.
2. Background
2.1. Congestion Signals
In this section we are going over several congestion signals that REPS can take advantage of if they are implemented in the network. At minimum we require ECN to be supported by the switches.
- Delay:
End-to-end delay serves as a reliable indicator of congestion. Delay can be calculated by the sender through measuring the RTT, as seen in protocols like Swift (Kumar et al., 2020) and TCP (Cerf and Kahn, 1974). Alternatively, the receiver can annotate acknowledgment (ACK) packets with one-way delay information, a method used in protocols such as TIMELY (Mittal et al., 2015) and DX (Lee et al., 2017). In this context, we assume that precise delay measurements are obtained via NIC timestamping (Kumar et al., 2020; Bai et al., 2023).
- ECN-marking:
Explicit Congestion Notification (ECN) allows switches to indicate congestion by setting a bit in the traffic class field of the IP header when congestion is detected. This marked packet is then returned to the sender by the receiver, prompting the sender to adjust its load balancing or its transmission rate accordingly (Zhu et al., 2015; Alizadeh et al., 2010; Ye et al., 2019). Unlike time-based signals that provide detailed queuing delay information, ECN uses a single-bit mark, offering less granularity. Switches can employ various strategies for marking packets. For instance, in Random Early Detection (RED)(Floyd and Jacobson, 1993), switches probabilistically mark packets based on queue size, with the marking probability increasing linearly between two thresholds ( and ). Although ECN was originally intended to mark packets upon enqueue(Floyd et al., 2001), marking at dequeue has been shown to enable quicker response times for congestion control (CC) algorithms. This dequeue marking method is easily implemented on most existing switches (Wu et al., 2012). For this reason, the rest of the paper assumes the use of RED with dequeue marking for ECN packets. While ECN-based CC algorithms can respond faster than those based on delay, they may struggle with incast scenarios and can be difficult to fine-tune (Kumar et al., 2020). Finally, while delay accumulates over multiple hops, ECN would be able to filter out such cases where there is not really a significant congestion happening but only a small amount of queueing over multiple hops.
- Packet Losses:
Packet loss has long been a key indicator of severe congestion in networks (Floyd and Jacobson, 1993; Nichols and Jacobson, 2012). However, using packet losses as the sole signal for congestion detection can result in delayed responses, as losses typically indicate a point of significant congestion. Additionally, packet loss detection often relies on timeouts, which can be challenging to calibrate and may lead to unnecessary retransmissions. Having said that, timeout can still be very useful in the case of network failures, especially when having out-of-order delivery support since we can not easily rely on a packet arriving out of order to detect losses.
- Packet Trimming:
Packet trimming is a technique where a switch removes portions of a packet, such as the payload, instead of discarding the entire packet. This method preserves crucial information, like headers, which enables the host to quickly identify and respond to congestion (Cheng et al., 2014; Handley et al., 2017; Olteanu et al., 2022). Researchers have shown that packet trimming can be implemented on switches such as Intel Tofino, Broadcom Trident 4, and NVidia Spectrum 2 by utilizing their capability to reroute packets to an alternate port when the primary egress queue is full (Adrian et al., 2022). This can be particularly helpful in detecting congestion events versus losses events, as the former would produce a trim even while the latter a timeout (excluding edge cases where even a trim would fail).
REPS utilizes ECN as its primary congestion signal to determine if a path is congested. Notably, ECN marking is performed at the egress to ensure that the most current information is captured. If packet trimming is supported, REPS leverages this feature to enhance its ability to distinguish between packets lost due to congestion and those lost due to network failures. In situations where packet trimming is not supported, REPS employs a straightforward timeout mechanism to detect packet loss within the network.
3. Recycled Entropy Packet Spray (REPS)
REPS is a simple load balancer scheme that does not need any switch support besides ECMP (Hopps, 2009). With ECMP, each switch computes a hash function over some packet header fields to determine which path to forward the packet on. One of these fields is the entropy (e.g., the source port or the IPv6 Flow Label), which REPS changes on a per-packet basis.
The key idea behind REPS is very simple: we want to explore new possible paths when congestion is detected while re-using paths where there is little to no congestion. This has the immediate advantage of being both simple and lightway since we can potentially store only the last good entropy that we have seen and which is to be reused. However, the slightly more advanced version of it uses a circular buffer of a fixed size where we save all the cached entropies.
For the first of packets, REPS explores all possible entropies. This is necessary since initially we do not have any information about the status of the network. The receiver copies the entropies from the received packet to the ACK, thus forwarding them back to the sender. Whenever an ACK arrives, if the ACK is not ECN-marked, the entropy carried by the ACK is cached. Intuitively, because the corresponding data packet did not experience congestion, we can send other packets on the same path, and REPS will thus reuse it for subsequent packets. Otherwise, if the ACK was ECN-marked, REPS selects the next entropy. This can be further extended by defining additional policies (e.g., excluding bad entropies after some time or storing entropies in a priority queue based on their RTTs and/or ECN marking), but we do not explore this further in this initial version of this work. The circular buffer design guarantees that even when receiving back-to-back ACKs with good entropies, these can be correctly stored and used. On the sending side, if we have multiple elements in the buffer, we will always first select the oldest entropy to make sure to not get stuck on old information.
We evaluate REPS’ advantages versus traditional packet spraying and ECMP in Sec. 4.2. Algorithm 1 describes the REPS pseudocode, and Fig. 1(a) shows an example of REPS’ behavior.
3.0.1. Interaction between CC and load balancing
We design REPS with the idea of it being capable of working with any transport layer that does not require in-order delivery of packets. Having said that, co-designing the congestion control protocol with the load balancer algorithm can bring several benefits. For these reasons, we optimize REPS to work with SMaRTT (Bonato et al., 2024). REPS effectively mitigates the effects of fabric congestion, such as those caused by ECMP hashing collisions, without necessitating a reduction in the congestion window size. This capability is particularly important in permutation workloads on non-oversubscribed networks, where proper traffic balancing can theoretically prevent congestion. To enable REPS to respond to such congestion, SMaRTT is designed to ignore individual ECN-marked packets unless multiple marked packets are received. Specifically, SMaRTT maintains an exponentially weighted moving average of ECN-marked packets, updating the variable using the formula . The congestion window remains unchanged as long as , meaning that the system does not react to congestion unless 25 percent or more of the acknowledgments indicate congestion—a threshold high enough to exclude transient load imbalances. This mechanism, referred to as Wait to Decrease (WTD), is illustrated in Fig. 1, which demonstrates how it significantly improves the flow completion time (FCT) for permutation workloads running on a 1,024-node non-oversubscribed fat tree network.
4. Evaluation
Simulation Setup
We ran simulations using and extending the htsim packet-level network simulator (Handley et al., 2017). Our simulations consider different fat-tree topologies with 1024 nodes and different level of oversubsciption ranging from zero to 8:1. This is relevant due to the relevance of such topologies in production datacenters (Singh et al., 2015). We configure the network with a 4 KiB MTU size, a bandwidth of 100 Gbps, and a switch traversal latency of 500 ns, reflecting the specifications of current-generation switches (Broadcom, 2024; De Sensi et al., 2020). To simplify the analysis, we assume uniform link lengths and latencies, with each link exhibiting a latency of 500 ns.
State-of-the-art Comparison
We pair REPS together with SMaRTT and compare it against packet spraying and ECMP. We use SMaRTT for all of them in order to have the same baseline regarding congestion control. We enable trimming for all the algorithms.
4.1. Non-Oversubscribed Networks
In this section we evaluate the performance of REPS in a simple setting where there is no ovserubscription and no failures, meaning the network is perfectly symmetrical. Intuitively, this is the best situation for packet spraying since evenly splitting the packets across multiple links should result in the best performance. However, as we can see in Fig.2(a), this is not the case as REPS still offers a 10% advantage over oblivious packet spraying. This is because of ECMP collisions that still happen with spraying. While over long period of time, each link will be evenly used, there will still be short term collisions happening that also increase the RTT of packets as seen in Fig.2(b).
4.2. Oversubscribed Networks
In Fig.3(a), we evaluate the performance of SMaRTT with various load balancing algorithms: REPS, ECMP(Hopps, 2009), and oblivious packet spraying (Dixit et al., 2013). We simulate a 4 MiB permutation workload on a 8:1 oversubscribed fat tree network. The results show that REPS achieves again approximately a 10% improvement in completion time over oblivious packet spraying due to its superior fairness. In contrast, per-flow ECMP results in job completion times that are 1.5 times higher than those achieved with packet spraying and REPS, primarily because of increased flow collisions (which are not depicted in the plot). However, the individual RTTs for per-flow ECMP are relatively low since SMaRTT reduces the congestion window size upon detecting congestion caused by hash collisions. Additionally, an analysis of the packet RTT distribution in Fig. 3(b) reveals that REPS maintains RTTs that are closer to the target with lower variability, indicating improved network utilization.
4.3. Failures
We now evaluate what happens in a network where some links are failing, in particular we fail a fixed number of links between the aggregation switches and core switches of the fat tree. REPS now gains a 50% advantage (Fig. 4(b)) over spraying because it is able to avoid more efficiently the failing links, while spraying would randomly select failing paths over time and hence increase the completion time. For ECMP, certain flows don’t properly terminate as they are not able to adaptively avoid failing links.
5. Conclusion
We design REPS, a simple and resource-efficient load balancer that can be paired with different congestion control algorithms. REPS significantly improves the FCT, runtime and number of dropped packets as shown in our evaluation. REPS can drastically improve the load balancing performance compared to traditional per-flow ECMP or oblivious packet spraying at minimal complexity cost. We also show that REPS can effectively handle faults and network asymmetries. In a future version of this work we will expand REPS to be used alongside several congestion control algorithms and with different failure modes.